# Dynamical and structural properties of an absorbing phase transition: a case study from granular systems

### v1

@@ Abstract
We investigate the dynamical and structural properties of absorbing phase transitions (APTs) within granular systems. Specifically, we examine a model for vibrofluidized systems of spherical grains, which transition from a state of pure vertical motion to a horizontal diffusive state upon increased density. Depending on the details of the system, we observe both continuous and discontinuous transitions in numerical simulations. We explain this using a theoretical analysis based on kinetic theory applied to a coarse-grained model, which elucidates the role of a synchronization effect in determining the nature of the transition. A fluctuating hydrodynamic theory, which quantitatively describes the structural and dynamical properties of the active state such as hyperuniformity is derived from the microscopic dynamics, together with an equilibrium-like assumption concerning the noises on the hydrodynamic fields. This work expands on previous studies by providing a comprehensive examination of the APT characteristics and proposing new theoretical models to interpret the observed behaviors.
@@ I Introduction
In the context of non-equilibrium statistical physics, an absorbing phase occurs when a dynamical system is trapped in a stationary or periodic state in which no more evolution takes place. Upon varying a suitable control parameter, a transition to a diffusive state can occur. Such transitions are termed absorbing phase transitions (APTs). The characteristics and universality class of these transitions have been extensively discussed in the literature [1, 2]. APTs have been observed in many different contexts, from epidemiology [3] to sandpile models [4], and from quantum systems [5, 6, 7, 8, 9] to chemical reactions [10]. Since the pioneering work by Pine and coworkers [11], particle systems have also become a major area of study for APTs. Specifically, APTs have been observed in low Reynolds number reversible suspensions [12, 11, 13, 14, 15, 16], dense colloidal systems [17, 18, 19, 20], emulsions [21, 22], liquid crystals [23], glasses [24, 25, 26], jammed systems [27, 28, 29], granular materials [30, 31], chemical systems [32, 33, 34], ecology [35, 36, 37, 38] materials at yielding [39] and turbulent systems [40, 41, 42, 43, 44].
@@
In several of the cases mentioned above [45, 11, 13, 14, 23, 29, 41], the transition manifests in a continuous manner. This behavior has been interpreted within the framework of directed percolation  [46] or conserved directed percolation models, notably the Manna models [47, 48]. These theoretical approaches have proven valuable in understanding the nature of the continuous transitions observed. However, a contrasting phenomenon has been documented in many other situations [49, 15, 50, 51]. In these instances, researchers have found that the transition exhibits a discontinuous character. To explain this divergence from the continuous behavior, it has been speculated that certain interactions might play a crucial role. Specifically, interactions potentially arising from hydrodynamic effects or material elasticity have been suggested as possible mechanisms responsible for the discontinuous nature of the transition [52].
@@
Recently, we introduced a simple vibrofluidized granular model in which macroscopic spherical beads, confined within a quasi-2D cell, undergo a transition from a state in which they are locked in a vertical motion to a diffusive and active state in the horizontal direction as the density is increased [53]. Furthermore, by adjusting parameters like the confinement height or the vibration amplitude, we showed that we can induce either a continuous or discontinuous transition. While continuous to discontinuous absorbing transitions have been observed in various systems [54, 55, 56, 57, 58, 59, 52, 35, 9], the underlying physical explanations have often remained unclear. In our case, we demonstrated that this phenomenon was related to the synchronization between the vibrating cell and the vertical motion of the beads. The granular system under study was described with two different models: a realistic one which can be numerically integrated through the discrete element method and a collisional coarse-grained one which can be simulated through event-driven molecular dynamics and provides a good starting point for analytical investigations based on kinetic theory and hydrodynamics.
@@
In this paper, we expand the scope of our investigation in two primary directions. First, we examine the dynamical and structural properties of the absorbing phase transition, aiming to elucidate potential distinctions between continuous and discontinuous cases. Our analysis of the absorbing phase transition’s properties focuses on identifying characteristic differences between continuous and discontinuous transitions. This examination of dynamical and structural aspects aims to deepen our understanding of the underlying mechanisms driving these transitions. Second, we explore theoretical models that can interpret the observed phase transition by extending the kinetic theory framework we proposed in Ref. 53 and by proposing an approach grounded in fluctuating hydrodynamics to study the active state. These complementary theoretical frameworks offer valuable perspectives on the complex dynamics of the system under study.
@@ II Absorbing phase transitions in a granular system
In this section, we describe two different numerical approaches used in the present study. First, we consider a realistic model, which is based on a detailed and accurate description of grain interactions and which includes the complete quasi-2D dynamics of the quasi-2D granular system. The advantage of this approach is that it models grain collisions in a more physically accurate manner, incorporating the mechanical properties of the materials involved. However, this approach turns out to be computationally very demanding. The second model we present is a 2D coarse-grained model that we have developed to simplify the problem [53]. This approach offers two significant benefits. These simulations become significantly less computationally demanding, enabling the exploration of time scales and system sizes increased by three orders of magnitude. Additionally, while our simplified model is approximate, it provides a better understanding of the underlying physical principles behind the observed phenomena and represent a more viable starting for the analytical approach discussed later in the paper.
II.1 The realistic model
Our first model consists of a vertically vibrating granular system confined between two plates in a quasi-2D geometry (see Fig. 1). The evolution is simulated using molecular dynamics simulations based on discrete element method [60, 61] using LAMMPS [62, 63]. This numerical approach is widely adopted to model realistic in-silico setups [64, 65], utilizing precise contact mechanics models [66, 67, 61, 68] for grain-grain and grain-plate interactions. These models encompass normal and tangential forces, capturing both their elastic and dissipative aspects. In particular, in this work, we adopt the Hertz-Mindlin contact model [69, 70]. Particles are also subject to a constant downward gravity field $g$.
@@ II.1 The realistic model
Within the LAMMPS simulations, these parameters are implemented in SI units, and we refer to Ref. 53 and its Supplementary Informations for the numerical values of the physical properties of the beads and additional information concerning the simulations details.
@@
Figure 1: Numerical quasi-2d geometry used in realistic simulations. A vertical displacement $z_{p}(t)$ is imposed to the box in order to provide external energy to the system. Because of tangential frictional forces, the grains lose horizontal energy during collisions with the top and bottom walls. This mechanism introduces an effective dissipation rate $\bar{\gamma}$ for the horizontal dynamics. Energy transfer between $z$ and $xy$ directions occurs during grain-grain non-planar collisions. This introduces an effective energy gain at collision for the $xy$-motion.
We consider a system of $N$ particles initialized with random velocities. At low densities, collisions are not frequent enough to convert the continuous supply of vertical energy into horizontal one. Consequently, the system gradually cools due to friction with the plates until it becomes completely arrested in the $xy$ plane, while still bouncing up and down along the $z$ axis. We refer to this state as the absorbing state, as the system cannot escape the $xy$ spatial configurations it has fallen into. At intermediate densities, the system reaches a horizontally active state in the $xy$ plane, where the energy “injected” during collisions is sufficient to balance the energy dissipated through tangential collisions with the plates. At very high densities, the system start to crystallize [71, 72, 73, 74, 75] into an hexagonal phase and reaches again an “absorbing phase” with 0 diffusivity in the $xy$ plane, consistent with previous studies on absorbing phase transitions in cyclically sheared suspensions [29, 13, 11, 29].
[…]
@@
[…]
The behavior of $T^{ss}$, is also presented in Fig. 2. Its value remains zero up to the absorbing-to-active transition at $\phi_{c}$. Following this transition, the order parameter monotonically increases, which is in contrast to the non-monotonic trend observed for diffusivity due to the spatial localization of the particles at higher density [29]. The choice of $T^{ss}$ as the order parameter will prove particularly suitable for the theoretical treatment we propose in Sec. IV.
@@
Interestingly, as discussed in Ref. 53, the nature of the transition as the packing fraction varies can be either continuous or discontinuous, depending on the values of physical parameters such as the amplitude of shaking or the height of the plate. This phenomenon was explained by an additional dynamical effect related to the synchronization of the beads with the moving plates. In what follows, we review and discuss this mechanism in detail.
From numerical simulations, we observe that an isolated particle, for a given set of parameters, can either begin oscillating in the $z$ direction in phase with the plate (a state we call synchronized) after a characteristic time $\tau_{s}$, or remain in chaotic motion, not synchronized with the plate. This crossover between periodic and chaotic dynamics for bouncing grains colliding with moving walls has been already observed in different experiments [76, 77, 78, 79, 80, 81] and understood theoretically by means of simple dynamical models [82, 83, 84]. In order to quantify the degree of synchronization, we use the following parameter [85]:
@@
It is important to note that synchronization of the particles requires a typical time $\tau_{s}$, which plays a key role in determining the critical packing fraction $\phi_{c}$ where the transition takes place. This timescale can be tuned via the shaking amplitude and the sample height. To illustrate this, we plot in Fig. 3b the behavior of $\phi_{c}$ as a function of $\tau_{s}$ for a range of different conditions. The critical packing fraction grows with the synchronization frequency $1/\tau_{s}$. As in Ref. 53, we measure $\tau_{s}$ for every $A$ and $h$ using an exponential fit of $s(t)$ starting from an asynchronized and absorbing state.
@@
Figure 3: Comparison of the realistic model (a) and b)) with the coarse-grained model (c) and d)). a) Evolution of the $xy$ temperature as a function of the density for various amplitude and height in the realistic model. The values of these parameters for each curve are indicated on the inset with corresponding markers. The inset is the synchronization map of the absorbing state with varying $h$ and $A$. When the absorbing state is synchronized, the transition is discontinuous, while the opposite is observed when the absorbing state is chaotic. b) Critical packing fraction as a function of $1/\tau_{s}$ in the realistic model. The synchronization time controls the packing fraction of the transition. c) Temperature as a function of the packing fraction in the coarse-grained model. Without synchronization ($01/\tau_{s}=$ 0), the transition is continuous, while at finite $\tau_{s}$, the transition is discontinuous (see however Sec. IV.2 for a discussion of the existence of a tricritical point at finite $\tau_{s}$). d) Critical packing fraction as a function of the synchronization time for various damping in the coarse-grained model. Consistently with the realistic mode, the synchronization time increases the critical packing fraction since it facilitates the transition.
@@
In summary, the observed phenomenology is attributable to the interplay between two distinct transitions: one between absorbing and active states in the $xy$-motion, and the other between synchronized and asynchronized states in the $z$-dynamics. Without synchronization, the system reaches continuously an absorbing state due to tangential friction with the plate while with synchronization, the transition is discontinuous and driven by purely dissipative collisions. Although it is possible to develop a theory for this system directly [86, 87, 88, 89, 90, 91], we instead recall, in the following subsection, the model introduced in Ref. 53, which offers a simplified framework for analyzing the phenomenology observed in the more realistic setting.
II.2 The coarse-grained model
@@
In the preceding section, we have accumulated a wealth of information from the realistic model. Here, we will introduce a simplified 2D coarse-grained model that incorporates the three primary mechanisms derived from the analysis of the realistic model. These mechanisms are: an effective dissipation rate for the horizontal motion, an effective parameter for the $z$-to-$xy$ energy transfer at grain-grain collisions and a synchronization time $\tau_{s}$. This simplified model will allow us to understand the interplay between APT and synchronization, and will rely on a minimal set of physical parameters compared to the realistic model.
@@
The beads are modeled as identical hard disks of mass $m$ and radius $r$ undergoing dissipative collisions characterized by a coefficient of restitution $\alpha$. In the quasi-2D realistic model, the energy is injected through vertical vibrations and subsequently converted into $xy$ momentum through off-plane collisions. Brito et al. [92] proposed a purely 2D model effectively replicating the energy transfer mechanism of the realistic model by introducing the following collision rule between particle $i$ and $j$:
@@
The last term in Eqs. (5) accounts for energy injection, with $\Delta_{ij}>0$ ensuring that energy input from this term remains positive at each collision, thus reproducing the conversion of vertical motion into horizontal motion in the realistic system. However, the collision can still be dissipative if the relative velocity is large. The $\Delta_{ij}$ is an effective parameter that takes into account the effects of the height of the plate and the driving on the energy transfer at collision on a coarse level. We will show that it can be chosen to depend on the degree of synchronization of the particles.
Furthermore, in the realistic model, the beads lose energy upon collisions with the ceiling or bottom of the plate. To incorporate this effect in our coarse-grained model, we introduce a viscous drag $\gamma$ during the free flight between collisions:
@@
For fixed parameters $\alpha$, $\gamma$, and $\Delta$, the system can exhibit either an absorbing state or an active state, depending on density. At low density, the dynamics are primarily governed by viscous drag, causing the system to eventually come to rest in an absorbing state. In contrast, at high density, the dynamics are mainly driven by collisions leading the system to an active steady state. It should be stressed that, up to this point, the dynamics we consider are very close to the ones introduced in Ref. 93.
To model the dynamics observed with the 3D realistic model, we must incorporate synchronization into our coarse-grained model as well. In the realistic model, simulations of free (non-colliding) particles with random initial conditions demonstrate a tendency to synchronize with the plate after a randomly distributed time interval $\tau_{s}$. Moreover, we recall that collisions between the particles disrupt their synchronization, as these collisions effectively randomize the post-collision vertical velocity. Even in-plane collisions between synchronized particles slightly asynchronize them over a timescale $\tau_{s}^{\prime}$ which is on average significantly shorter than $\tau_{s}$. To easily incorporate all these aspects of the realistic model into our model, we make the assumption that a single constant synchronization time, denoted as $\tau_{s}$, is sufficient to describe qualitatively the synchronization. In the coarse-grained model, particles are either labeled as synchronized or asynchronized. Free flight lasting more than $\tau_{s}$ synchronize the particle, while any collisions desynchronize it.
We recall that this synchronization effect plays a major role in the energy transfer at collision. In the realistic model, since synchronized particles colliding have the same $z$ coordinate, no transfer of vertical kinetic energy to the horizontal degrees of freedom happens. Asynchronized collisions are instead mostly non-planar, causing a significant $z$-to-$xy$ energy transfer. To capture this behavior in our existing coarse-grained model, we define a synchronization dependent $\Delta_{ij}$. If both particles involved in a collision have not had a collision for a time period longer than $\tau_{s}$, the collision does not add energy to the system, i.e. we set $\Delta_{ij}=0$. For all other collisions, where at least one of the particles involved has had a collision within the last $\tau_{s}$ time period, we process the collision with a fixed $\Delta_{ij}>0$, representing the transfer of energy and momentum resulting from asynchronized collisions.

### v2

@@ Abstract
We investigate the dynamical and structural properties of absorbing phase transitions (APTs) within granular systems. Specifically, we examine a model for vibrofluidized systems of spherical grains, which undergo a transition from a state of purely vertical motion to one characterized by horizontal diffusion as the density increases. Numerical simulations reveal that, depending on the specific system parameters, both continuous and discontinuous transitions can occur, each associated with markedly distinct structural properties at the transition point. We explain this using a theoretical analysis based on kinetic theory applied to an effective 2D model, which elucidates the role of a synchronization effect in determining the nature of the transition. A fluctuating hydrodynamic theory, which quantitatively describes the structural and dynamical properties of the active state such as hyperuniformity is derived from the microscopic dynamics, together with an equilibrium-like assumption concerning the noises on the hydrodynamic fields. This work expands on previous studies by providing a comprehensive examination of the APT characteristics and proposing new theoretical models to interpret the observed behaviors.
@@ I Introduction
In the context of non-equilibrium statistical physics, an absorbing phase occurs when a dynamical system is trapped in a stationary or periodic state in which no more evolution takes place. Upon varying a suitable control parameter, a transition to a diffusive state can occur. Such transitions are termed absorbing phase transitions (APTs). The characteristics and universality class of these transitions have been extensively discussed in the literature [hinrichsen2000non, lubeck2004universal]. APTs have been observed in many different contexts, from epidemiology [Mata2021] to sandpile models [Dickman1998, hinrichsen1999flowing, hinrichsen2000flowing], and from quantum systems [PhysRevResearch.3.013238, PhysRevB.109.L020304, PhysRevLett.130.120402, chertkov2023characterizing, marcuzzi2016absorbing] to chemical reactions [ziff1986kinetic]. Since the pioneering work by Pine and coworkers [pine2005chaos], particle systems have also become a major area of study for APTs. Specifically, APTs have been observed in low Reynolds number reversible suspensions [reichhardt2023reversible, pine2005chaos, corte2008random, nagamanasa2014experimental, jeanneret2014geometrically, franceschini2011transverse], dense colloidal systems [keim2014mechanical, tjhung2015hyperuniform, tjhung2016criticality, PhysRevLett.125.148001], emulsions [knowltonsm14, Weijs2015], liquid crystals [takeuchi2007directed], glasses [fiocco2013, regev2013onset, bhaumik2021role], jammed systems [PhysRevLett.131.238202, PhysRevLett.127.038002, Ness2020], granular materials [Mobius2014, neel2014dynamics], chemical systems [vellela2009stochastic, schlogl1972chemical, de2023effects], ecology [garcia2024interactions, pascual2005criticality, clar1997phase, chang2020modelling] materials at yielding [PhysRevLett.132.268203] and turbulent systems [bertin2024cascade, chantry2017universal, PhysRevLett.132.264002, hof2023directed, lemoult2016directed].
@@
In several of the cases mentioned above [jocteur2025random, pine2005chaos, corte2008random, nagamanasa2014experimental, takeuchi2007directed, Ness2020, chantry2017universal], the transition manifests in a continuous manner. This behavior has been interpreted within the framework of directed percolation  [Rossi2000] or conserved directed percolation models, notably the Manna models [LeDoussal2015, Martiniani2019]. These theoretical approaches have proven valuable in understanding the nature of the continuous transitions observed. However, a contrasting phenomenon has been documented in many other situations [Kawasaki2016, jeanneret2014geometrically, leishangthem2017yielding, di2016self]. In these instances, researchers have found that the transition exhibits a discontinuous character. To explain this divergence from the continuous behavior, it has been speculated that certain interactions might play a crucial role. Specifically, interactions potentially arising from hydrodynamic effects or material elasticity have been suggested as possible mechanisms responsible for the discontinuous nature of the transition [mari2022absorbing].
@@
Recently, we introduced a simple vibrofluidized granular model in which macroscopic spherical beads, confined within a quasi-2D cell, undergo a transition from a state in which they are locked in a vertical motion to a diffusive and active state in the horizontal direction as the density is increased [maire2024interplay]. Furthermore, by adjusting parameters like the confinement height or the vibration amplitude, we showed that we can induce either a continuous or discontinuous transition. While continuous to discontinuous absorbing transitions have been observed in various systems [Fisher97, van1998wilson, pires2023tricritical, assis2009discontinuous, PhysRevE.93.012110, Jagla2014, mari2022absorbing, garcia2024interactions, marcuzzi2016absorbing], the underlying physical explanations have often remained unclear. In our case, we demonstrated that this phenomenon was related to the synchronization between the vibrating cell and the vertical motion of the beads. The granular system under study was described with two different models: a realistic quasi-2D one which can be numerically integrated through the discrete element method and an effective one in 2D which can be simulated through event-driven molecular dynamics and provides a good starting point for analytical investigations based on kinetic theory and hydrodynamics.
@@
In this paper, we expand the scope of our previous investigation [maire2024interplay] in two primary directions. First, we examine the dynamical and structural properties of the absorbing phase transition, aiming to elucidate potential distinctions between continuous and discontinuous cases. Our analysis of the absorbing phase transition’s properties focuses on identifying characteristic differences between continuous and discontinuous transitions. This examination of dynamical and structural aspects aims to deepen our understanding of the underlying mechanisms driving these transitions and to draw connections with analogous phenomena in other systems.. Second, we explore theoretical models that can interpret the observed phase transition by extending the kinetic theory framework we proposed in Ref. maire2024interplay and by proposing an approach grounded in fluctuating hydrodynamics to study the active state. These complementary theoretical frameworks offer valuable perspectives on the complex dynamics of the system under study.
@@ II Absorbing phase transitions in a granular system
In this section, we describe two different numerical approaches used in the present study. First, we consider a realistic quasi-2D model, which is based on a detailed and accurate description of grain interactions and which includes the complete quasi-2D dynamics of the quasi-2D granular system. The advantage of this approach is that it models grain collisions in a more physically accurate manner, incorporating the mechanical properties of the materials involved. However, this approach turns out to be computationally very demanding. The second model we present is an effective 2D model that we have developed to simplify the problem [maire2024interplay]. This approach offers two significant benefits. These simulations become significantly less computationally demanding, enabling the exploration of time scales and system sizes increased by three orders of magnitude. Additionally, while our simplified model is approximate, it provides a better understanding of the underlying physical principles behind the observed phenomena and represent a more viable starting for the analytical approach discussed later in the paper.
II.1 The realistic quasi-2D model
Our first model consists of a vertically vibrating granular system confined between two plates in a quasi-2D geometry (see Fig. 1). The evolution is simulated using molecular dynamics simulations based on discrete element method [Cundall79, PoeschelBook] using LAMMPS [Plimpton2022, kloss2010granular]. This numerical approach is widely adopted to model realistic in-silico setups [Plati2019, Osullivan2009], utilizing precise contact mechanics models [Zhang2005, Plati2022, PoeschelBook, PopovBook] for grain-grain and grain-plate interactions. These models encompass normal and tangential forces, capturing both their elastic and dissipative aspects. In particular, in this work, we adopt the Hertz-Mindlin contact model [mindlin1949compliance, LammpsSiteGran]. Particles are also subject to a constant downward gravity field $g$.
@@ II.1 The realistic quasi-2D model
Within the LAMMPS simulations, these parameters are implemented in SI units, and we refer to Ref. maire2024interplay and its Supplementary Informations for the numerical values of the physical properties of the beads and additional information concerning the simulations details.
@@
Figure 1: Numerical quasi-2D geometry used in realistic simulations. A vertical displacement $z_{p}(t)$ is imposed to the box in order to provide external energy to the system. Because of tangential frictional forces, the grains lose horizontal energy during collisions with the top and bottom walls. This mechanism introduces an effective dissipation rate $\bar{\gamma}$ for the horizontal dynamics. Energy transfer between $z$ and xyxy directions occurs during grain-grain non-planar collisions. This introduces an effective energy gain at collision for the xyxy-motion.
We consider a system of $N$ particles initialized with random velocities. At low densities, collisions are not frequent enough to convert the continuous supply of vertical energy into horizontal one. Consequently, the system gradually cools due to friction with the plates until it becomes completely arrested in the $xy$ plane, while still bouncing up and down along the $z$ axis. We refer to this state as the absorbing state, as the system cannot escape the $xy$ spatial configurations it has fallen into. At intermediate densities, the system reaches a horizontally active state in the $xy$ plane, where the energy “injected” during collisions is sufficient to balance the energy dissipated through tangential collisions with the plates. At very high densities, the system start to crystallize [Reis2006, Castillo2012, Castillo2015, Olafsen98, Komatsu2015] into an hexagonal phase and reaches again an “absorbing phase” with 0 diffusivity in the $xy$ plane, consistent with previous studies on absorbing phase transitions in cyclically sheared suspensions [Ness2020, corte2008random, pine2005chaos, Ness2020].
This qualitative picture is illustrated in Fig. 2 where the diffusivity in the $xy$ plane is plotted as a function of the area fraction $\phi=\pi N\sigma^{2}/4L^{2}$. The diffusivity is given by $D=\lim_{t\to\infty}\langle(\bm{r}_{xy}(t)-\bm{r}_{xy}(0))^{2}\rangle/4t$, where $\bm{r}_{xy}$ represents the horizontal coordinates of the position of a given particle, and $\langle\cdot\rangle$ refers to the average over all the grains. Indeed, at $\phi^{(1)}_{c}$, we observe the first transition, due to the competition between collision and dissipation, while at $\phi_{c}^{(2)}$ we observe the second, re-entrant transition driven by cage formation and crystallization. In this paper, we focus entirely on the first transition, whose critical packing fraction will be simply denoted as $\phi_{c}$.
@@
$T=\frac{1}{\mathcal{T}}\int_{t_{0}}^{t_{0}+\mathcal{T}}\frac{m}{2}\langle v_{x}^{2}(t)+v_{y}^{2}(t)\rangle dt$ (1)
where $\langle\cdot\rangle$ refers to the average over all the particles in an instantaneous configuration, $t_{0}$ is the initial measurement time and $\mathcal{T}$ is the observation time. This plays the role of a macroscopic effective temperature, sometimes called the granular temperature. We define $T_{ss}$ as the steady-state temperature, obtained by choosing $t_{0}$ to be larger than any relaxation time in the system.
The behavior of $T^{ss}$, is also presented in Fig. 2. Its value remains zero up to the absorbing-to-active transition at $\phi_{c}$. Following this transition, the order parameter monotonically increases, which is in contrast to the non-monotonic trend observed for diffusivity due to the spatial localization of the particles at higher density [Ness2020]. The choice of $T^{ss}$ as the order parameter will prove particularly suitable for the theoretical treatment we propose in Sec. IV.
@@
Interestingly, as discussed in Ref. maire2024interplay, the nature of the transition as the packing fraction varies can be either continuous or discontinuous, depending on the values of physical parameters such as the amplitude of shaking or the height of the plate. This phenomenon was explained by an additional dynamical effect related to the synchronization of the beads with the moving plates. In what follows, we review and discuss this mechanism in detail.
From numerical simulations, we observe that an isolated particle, for a given set of parameters, can either begin oscillating in the $z$ direction in phase with the plate (a state we call synchronized) after a characteristic time $\tau_{s}$, or remain in chaotic motion, not synchronized with the plate. This crossover between periodic and chaotic dynamics for bouncing grains colliding with moving walls has been already observed in different experiments [Pieraski1983, Pieraski1985, Chastaing2015, mujica2016dynamics, schindler2023phase, rivas2011sudden] and understood theoretically by means of simple dynamical models [everson1986chaotic, Metha1990, Luck1993]. In order to quantify the degree of synchronization, we use the following parameter [acebron2005kuramoto]:
@@
It is important to note that synchronization of the particles requires a typical time $\tau_{s}$, which plays a key role in determining the critical packing fraction $\phi_{c}$ where the transition takes place. This timescale can be tuned via the shaking amplitude and the sample height. To illustrate this, we plot in Fig. 3b the behavior of $\phi_{c}$ as a function of $(f\tau_{s})^{-1}$ for a range of different conditions. We find that the critical packing fraction measured for different amplitudes grows with the synchronization frequency $1/\tau_{s}$ at fixed $h$. As in Ref. maire2024interplay, we measure $\tau_{s}$ for every $A$ and $h$ using an exponential fit of $s(t)$ starting from an asynchronized and absorbing state.
@@
Figure 3: Comparison of the realistic quasi-2D model (a) and b)) with the effective 2D model (c) and d)). a) Evolution of the $xy$ temperature as a function of the density for various amplitude and height in the realistic quasi-2D model. The values of these parameters for each curve are indicated on the inset with corresponding markers. The inset is the synchronization map of the absorbing state with varying $h$ and $A$. When the absorbing state is synchronized, the transition is discontinuous, while the opposite is observed when the absorbing state is chaotic. b) Critical packing fraction as a function of $1/\tau_{s}$ in the realistic quasi-2D model. The synchronization time controls the packing fraction of the transition. c) Temperature as a function of the packing fraction in the effective 2D model. Without synchronization ($1/\tau_{s}=0$), the transition is continuous, while at finite $\tau_{s}$, the transition is discontinuous (see however Sec. IV.2 for a discussion of the existence of a tricritical point at finite $\tau_{s}$). d) Critical packing fraction as a function of the synchronization time for various damping in the effective 2D model. Consistently with the realistic model, the synchronization time increases the critical packing fraction since it facilitates the transition.
@@
In summary, the observed phenomenology is attributable to the interplay between two distinct transitions: one between absorbing and active states in the $xy$-motion, and the other between synchronized and asynchronized states in the $z$-dynamics. Without synchronization, the system reaches continuously an absorbing state due to tangential friction with the plate while with synchronization, the transition is discontinuous and driven by purely dissipative collisions. Although it is possible to develop a theory for this system directly [maynar2019understanding, maynar2019homogeneous, brey2019inhomogeneous, brey2020self, brey2020kinetic, mayo2023confined], we instead recall, in the following subsection, the model introduced in Ref. maire2024interplay, which offers a simplified framework for analyzing the phenomenology observed in the more realistic setting.
II.2 The effective 2D model
@@
In the preceding section, we have accumulated a wealth of information from the realistic quasi-2D model. Here, we will introduce an effective 2D model that incorporates the three primary mechanisms derived from the analysis of the realistic quasi-2D model. These mechanisms are: an effective dissipation rate for the horizontal motion, an effective parameter for the $z$-to-xyxy energy transfer at grain-grain collisions and a synchronization time $\tau_{s}$. This simplified model will allow us to understand the interplay between APT and synchronization, and will rely on a minimal set of physical parameters compared to the realistic quasi-2D model.
@@
The beads are modeled as identical hard disks of mass $m$ and radius $r$ undergoing dissipative collisions characterized by a coefficient of restitution $\alpha$. In the realistic quasi-2D model, the energy is injected through vertical vibrations and subsequently converted into $xy$ momentum through off-plane collisions. Brito et al. [brito2013hydrodynamic] proposed a purely 2D model effectively replicating the energy transfer mechanism of the realistic quasi-2D model by introducing the following collision rule between particle $i$ and $j$:
@@
The last term in Eqs. (5) accounts for energy injection, with $\Delta_{ij}>0$ ensuring that energy input from this term remains positive at each collision, thus reproducing the conversion of vertical motion into horizontal motion in the realistic quasi-2D system. However, the collision can still be dissipative if the relative velocity is large. The $\Delta_{ij}$ is an effective parameter that takes into account the effects of the height of the plate and the driving on the energy transfer at collision on a coarse level. We will show that it can be chosen to depend on the degree of synchronization of the particles.
Furthermore, in the realistic quasi-2D model, the beads lose energy upon collisions with the ceiling or bottom of the plate. To incorporate this effect in our effective 2D model, we introduce a viscous drag $\gamma$ during the free flight between collisions:
@@
For fixed parameters $\alpha$, $\gamma$, and $\Delta$, the system can exhibit either an absorbing state or an active state, depending on density. At low density, the dynamics are primarily governed by viscous drag, causing the system to eventually come to rest in an absorbing state. In contrast, at high density, the dynamics are mainly driven by collisions leading the system to an active steady state. It should be stressed that, up to this point, the dynamics we consider are very close to the ones introduced in Ref. lei2019hydrodynamics.
To model the dynamics observed with the realistic quasi-2D model, we must incorporate synchronization into our effective 2D model as well. In the realistic quasi-2D model, simulations of free (non-colliding) particles with random initial conditions demonstrate a tendency to synchronize with the plate after a randomly distributed time interval $\tau_{s}$. Moreover, we recall that collisions between the particles disrupt their synchronization, as these collisions effectively randomize the post-collision vertical velocity. Even in-plane collisions between synchronized particles slightly asynchronize them over a timescale $\tau_{s}^{\prime}$ which is on average significantly shorter than $\tau_{s}$. To easily incorporate all these aspects of the realistic quasi-2D model into our model, we make the assumption that a single constant synchronization time, denoted as $\tau_{s}$, is sufficient to describe qualitatively the synchronization. In the effective 2D model, particles are either labeled as synchronized or asynchronized. Free flight lasting more than $\tau_{s}$ synchronize the particle, while any collisions desynchronize it.
We recall that this synchronization effect plays a major role in the energy transfer at collision. In the realistic quasi-2D model, since synchronized particles colliding have the same $z$ coordinate, no transfer of vertical kinetic energy to the horizontal degrees of freedom happens. Asynchronized collisions are instead mostly non-planar, causing a significant $z$-to-xyxy energy transfer. To capture this behavior in our existing effective 2D model, we define a synchronization dependent $\Delta_{ij}$. If both particles involved in a collision have not had a collision for a time period longer than $\tau_{s}$, the collision does not add energy to the system, i.e. we set $\Delta_{ij}=0$. For all other collisions, where at least one of the particles involved has had a collision within the last $\tau_{s}$ time period, we process the collision with a fixed $\Delta_{ij}>0$, representing the transfer of energy and momentum resulting from asynchronized collisions.